PLEASE HELP ME IF ANYONE CAN
FOR WHICH OF THE FOLLOWING EQUATIONS IS X=-2 a ROOT?
(1)2/X-X=0
(2)X^2+4=0
(3)X^7+4X+4=0

Answer by CharlesG2(834)

PLEASE HELP ME IF ANYONE CAN
FOR WHICH OF THE FOLLOWING EQUATIONS IS X=-2 a ROOT?
(1)2/X-X=0
(2)X^2+4=0
(3)X^7+4X+4=0
did all 3 below
(1) 2/x - x = 0
(2 - x^2)/x = 0 (using common denominator x)
2 - x^2 = 0
-x^2 = -2
x^2 = 2
((( x = +- sqrt(2) )))
(2) x^2 + 4 = 0
x^2 = -4
x^2 = -1 * 4
x = i * +-2
x = +- 2i (i is the imaginary root of -1)
(3) x^7 + 4x + 4 = 0
well I certainly not gonna do this one
I think you mean x^2 + 4x + 4 = 0 and did a typing error
x^2 + 4x + 4 = 0
(x + 2)(x + 2) = 0 (use FOIL (first outer inner last) and you will see it equals previous line)
x = -2